maths89898:

A soda company would like to replace the label on a can of soda with a radius of 3 and one fourth cm. If the label touches end to end with no overlap, what is the length of the label? Use 22 over 7 for π. 143 over 56 cm 143 over 28 cm 143 over 14 cm 143 over 7 cm

CHICKENNOODLES:

the circumference has to be equal to the total length of its boundaries

CHICKENNOODLES:

and the length of the label is 143/7cm so the answer would be d

maths89898:

wait wait howd u work that out tho

maths89898:

i trust ur answer but i wanna know how to solve this on my own

CHICKENNOODLES:

so the radius of the circle r=3+1/4 =13/4cm C = 2π × r => C = 2 × 22 / 7 × 13 / 4 => C = ( 11 × 13 ) / 7 => C = 143 / 7.

maths89898:

one more last question but howd u get 13/4cm from the radius?

CHICKENNOODLES:

The length of label is equivalent to the circumference of circle having r = 13 / 4 cm.

maths89898:

ty!

CHICKENNOODLES:

yw